# The computational power of discrete chemical reaction networks with bounded executions

### v3

@@ 6.3 Impossibility of stably deciding not eventually constant predicates
Assume WLOG that $\phi(\mathbf{x}_{0})=0$. For each $i\in\mathbb{N}$, let $\mathbf{v}_{i}=\mathbf{x}_{i+1}-\mathbf{x}_{i}$, noting by condition (2) that $\mathbf{v}_{i}\geq 0$.
We consider the sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\mathbf{a}_{2},\dots$ defined as follows. Let $\mathbf{a}_{0}$ be a stable configuration reachable from $\mathbf{x}_{0}$; since the correct answer is no, all species present in $\mathbf{a}_{0}$ vote no. Now add $\mathbf{v}_{0}$ to $\mathbf{a}_{0}$. By additivity, the configuration $\mathbf{a}_{0}+\mathbf{v}_{0}$ is reachable from $\mathbf{x}_{1}=\mathbf{x}_{0}+\mathbf{v}_{0}$. Since the correct answer for $\mathbf{x}_{1}$ is yes, $\mathcal{D}$ must go from $\mathbf{a}_{0}+\mathbf{v}_{0}$ to a stable “yes” configuration, call this $\mathbf{a}_{1}$. Now add $\mathbf{v}_{1}$ to $\mathbf{a}_{1}$. Since the correct answer is no, $\mathcal{D}$ must now reach from $\mathbf{a}_{1}+\mathbf{v}_{1}$ to a stable “no” configuration, call it $\mathbf{a}_{2}$. By condition (3), each $\mathbf{v}_{i}$ obeys $\|\mathbf{v}_{i}\|<c$ for some constant $c$.
@@
Since $\mathcal{D}$ is entirely execution bounded, by Theorem 6.5, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{w}\cdot\mathbf{x}$, where $\mathbf{w}\geq\mathbf{0}$. Adding $\mathbf{v}_{i}$ to $\mathbf{a}_{i}$ increases $\Phi$ by $\mathbf{w}(\mathbf{v}_{i})$, which is bounded above by a constant since $\|\mathbf{v}_{i}\|<c$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $\mathbf{v}_{i}$ and then going from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{w}(\mathbf{v}_{i})+2$).
@@
The statement of Theorem 6.9 does not mention the concept of a leader, but it would typically apply to leaderless CRDs. A CRD may be execution bounded from configurations with a single leader, but not execution bounded when multiple leaders are present (preventing the use of Theorem 6.5, which requires the CRD to be execution bounded from all configurations). For example, in Lemma 4.5, reaction (9) occurs finitely many times if the leader/voter $S_{Y}$ or $S_{N}$ has count 1. However, if $S_{Y}$ and $S_{N}$ can be present simultaneously (e.g., if we start with two leaders), then the reactions $S_{Y}+V_{NN}\mathop{\rightarrow}\limits S_{Y}+V_{YN}$ and $S_{N}+V_{YN}\mathop{\rightarrow}\limits S_{N}+V_{NN}$ can flip between $V_{NN}$ and $V_{YN}$ infinitely often in an unbounded execution.
@@
{lemmarep}
If a leaderless CRD or CRC is execution bounded, then it is entirely execution bounded.
[…]
submissionA proof is in the appendix. Since $\mathcal{C}$ is leaderless, the sum of two valid initial configurations is also valid. Thus if we can produce some species from a valid initial configuration, we can produce arbitrarily large counts of all species by adding up sufficiently many initial configurations. This means that for any configuration $\mathbf{x}$, from any sufficiently large valid initial configuration $\mathbf{i}$, some $\mathbf{y}\geqq\mathbf{x}$ is reachable from $\mathbf{i}$. But if $\mathcal{C}$ is execution bounded from $\mathbf{i}$, since $\mathbf{i}\Rightarrow\mathbf{y}$, it must also be execution bounded from $\mathbf{y}$, thus also from $\mathbf{x}$ since by additivity any reactions applicable to $\mathbf{x}$ are also applicable to $\mathbf{y}$.
Proof 6.23.
@@
Section 6.3 lets us replace “entirely execution bounded” in Section 6.3 with “leaderless and execution bounded”:
@@
In particular, since the original model of population protocols [1] defined them as leaderless and all-voting—and since population protocols are non-collapsing—we have the following.
@@
{toappendix}
6.4 Feedforward CRNs
We show that another common constraint, feedforwardness, significantly reduces computational power, making it impossible to decide even simple mod and threshold sets.
Definition 6.26.
A CRN is reaction-feedforward if reactions can be ordered $r_{1},r_{2},\ldots,r_{n}$ such that, for all $k<\ell$, no reactant of $r_{k}$ appears in $r_{\ell}$ (as either reactant or product).
Reaction-feedforward CRNs are significant in the sense that many continuous real-valued CRNs computing numerical-valued functions (where the count of some species $Y$ is interpreted as the output, e.g., $2X\to Y$ computes $f(x)=\lfloor x/2\rfloor$) can be computed by reaction-feedforward CRNs [4].666 The definition of feedforward in reference [4] is different from the definition given here, being based on an ordering of species rather than reactions. However, it is straightforward to verify by inspection that the CRNs given for the positive results of [4] are reaction-feedforward according to Definition 6.26. Compared to general CRNs, reaction-feedforward CRNs are easy to analyze and prove correctness. One reason is that, if a reaction-feedforward CRN can reach terminal configuration from $\mathbf{x}$ at all, then it is execution bounded from $\mathbf{x}$.
There is a similar definition, called simply feedforward in [4], based on ordering of species rather than reactions. We use the term species-feedforward to avoid confusion with Definition 6.26. We say a reaction $(\mathbf{r},\mathbf{p})$ produces a species $S$ if $\mathbf{p}(S)>\mathbf{r}(S)$, and it consumes $S$ if $\mathbf{r}(S)>\mathbf{p}(S)$.
Definition 6.27.
A CRN is species-feedforward if species can be ordered $S_{1},S_{2},\ldots,S_{n}$ such that every reaction producing a species $S_{\ell}$ consumes a earlier species $S_{k}$ where $k<\ell$.
Although the term “linear potential function” was not used in [4], it is shown in [4, Lemma 4.8] that species-feedforward CRNs have a linear potential function (assigning weight $\frac{1}{K^{i}}$ to species $S_{i}$ for a suitably large constant $K$), thus are entirely execution bounded. The same is not always true of reaction-feedforward CRNs, for example $X\mathop{\rightarrow}\limits 2X$ is reaction-feedforward but not execution bounded. However, we can use similar techniques to proofs used for so-called noncompetitive CRNs in [18] to show “reasonable” reaction-feedforward CRNs are execution bounded.
Lemma 6.28.
Suppose in a reaction-feedforward CRN that $\mathbf{i}\Rightarrow\mathbf{c}$ by execution $P$, and $\mathbf{i}\Rightarrow\mathbf{d}$ by execution $Q$. If any reaction occurs less in $P$ than $Q$, then $\mathbf{c}$ is not terminal.
Proof 6.29.
Here we equivalently think of an execution from $\mathbf{i}$ as a sequence of reactions, since from those and $\mathbf{i}$ we can deduce the configurations in the execution. Define #$\#(r_{k},P)$ as the number of times reaction $r_{k}$ occurs in the execution $P$. Let $r_{k}$ be the first reaction in the reaction-feedforward order such that #$\#(r_{k},P)<\#(r_{k},Q)$. Assume, for brevity of explanation, that $r_{k}$ has only one reactant, denoted $A$; the argument below, however, is general and applies to any number of reactants in $r_{k}$.
By the definition of a reaction-feedforward CRN, the reactions $r_{k+1}$ through $r_{n}$ do not affect the count of $A$. Further, reactions $r_{1}$ through $r_{k-1}$ can only produce $A$ and not consume it, reactions $r_{1}$ through $r_{k}$ can increase the count of $A$, and among them, only $r_{k}$ can decrease it. Let $m=\#(r_{k},P)$. Let $Q^{\prime}$ represent the prefix sequence $(\mathbf{i},\mathbf{x}_{1},\ldots,\mathbf{x}_{p})$ of $Q$ where the transition $\mathbf{x}_{p}\Rightarrow\mathbf{x}_{p+1}$ corresponds to the $(m+1)$st execution of reaction $r_{k}$. The configuration $\mathbf{x}_{p}$ is thus the configuration just before $r_{k}$ occurs more in $Q$ than in $P$.
Note that reactions $r_{1}$ through $r_{k-1}$ occur at least as often in $P$ as in $Q$ (i.e. #$\#(r_{i},P)\geq\#(r_{i},Q)$ for $i=1$ to $k-1$). Therefore, they occur at least as often in $P$ as in $Q^{\prime}$, since $Q^{\prime}$ is a prefix of $Q$. Moreover, by our choice of $Q^{\prime}$, #$\#(r_{k},P)=\#(r_{k},Q^{\prime})$. So $A$ is present in $\mathbf{c}$, i.e. $\mathbf{c}(A)>0$. Thus, $r_{k}$ is applicable at $\mathbf{c}$, so $\mathbf{c}$ is not terminal.
The following corollary implies that any reaction-feedforward CRN that can reach a terminal configuration from $\mathbf{i}$ is execution bounded from $\mathbf{i}$.
Corollary 6.30.
In a reaction-feedforward CRN $\mathcal{C}$, if there is a terminal configuration $\mathbf{c}_{\mathbf{i}}$ reachable from initial configuration $\mathbf{i}$, then $\mathbf{c}_{\mathbf{i}}$ is reached by every sufficiently long execution from $\mathbf{i}$, so $\mathbf{c}_{\mathbf{i}}$ is the only terminal configuration reachable from $\mathbf{i}$. Furthermore, all of these executions are permutations of the same number of each reaction type. In particular, $\mathcal{C}$ is execution bounded from $\mathbf{i}$.
Proof 6.31.
Let $P$ be the execution leading from $\mathbf{i}$ to $\mathbf{c}_{\mathbf{i}}$. Let $Q$ be any execution from $\mathbf{i}$.
We first claim that $|Q|\leq|P|$; suppose otherwise the sake of contradiction. By the pigeonhole principle, $Q$ must have more of some reaction than $P$ does. By Lemma 6.28, this would imply that $\mathbf{c}_{\mathbf{i}}$ is not terminal, a contradiction. Therefore, no execution $Q$ can be longer than $P$; in particular $\mathcal{C}$ is execution bounded from $\mathbf{i}$.
If $|Q|=|P|$, we claim $Q$ must be a permutation of $P$. Supposing otherwise, by the pigeonhole principle, $Q$ would have more of some reaction than $P$, again with Lemma 6.28 implying that $\mathbf{c}_{\mathbf{i}}$ is not terminal, a contradiction.
Finally, suppose $|Q|<|P|$. By pigeonhole some reaction occurs more frequently in $P$ than in $Q$, so by Lemma 6.28, $Q$ cannot reach a terminal configuration. Therefore all executions from $\mathbf{i}$ to $\mathbf{c}_{\mathbf{i}}$ are permutations of $P$.
As noted, in the model of continuous CRNs, it is known that all the functions that can be stably computed (the continuous, piecewise linear functions) can be stably computed by reaction-feedforward CRNs [4]. In contrast, with discrete CRNs computing predicates, we show that reaction-feedforward CRNs cannot stably decide all semilinear sets by giving two counterexamples, showing that reaction-feedforward CRDs can decide neither “the simplest” mod set (6.31) nor “the simplest” threshold set (6.32). Specifically, we chose the parity and majority predicate as our counterexamples, although the techniques generalize to more complex mod and threshold sets, e.g., $[X_{1}+2X_{2}\equiv 3\mod 5?]$.
{lemmarep}
Reaction-feedforward CRDs can’t stably decide the parity predicate $[X\equiv 1\mod 2?]$.
Proof 6.32.
We show that in any possible construction, the input species must be a reactant of two distinct reactions. By letting the CRN stabilize and then introducing another input molecule, there must exist a set of rules inverting the output in either way, consisting of at least two reactions with $X$ as reactant, breaking the reaction-feedforward condition.
Consider the set of even numbers. A simple, non-reaction-feedforward CRD that decides parity is:
$Y+X\rightarrow N$
$N+X\rightarrow Y$
where $X$ is the input species, $Y$ is a yes voter, and $N$ is a no voter, initialized with $1Y$ and $nX$. In either way to order these reactions, a reactant of the first reaction appears in the second reaction. Thus, the CRN is not reaction-feedforward.
To show that no such CRN could decide parity, we show that any construction requires us to have at least one reactant reappear in a later reaction, or even stronger: at least one species must be a reactant of two distinct reactions. Specifically, this is true for the input species $X$.
To motivate the choice of species, let’s consider an even simpler parity computing CRD.
$X+X\rightarrow Y$
$Y+X\rightarrow X$
where $X$ is both input and votes no, $Y$ votes yes, initialized with $1Y$ and $nX$. Only the input species appears twice as a reactant. Intuitively, this is true for all CRDs because we expect the input to be able to change our answer in either way, reversing the previous one.
Suppose for the sake of contradiction that there is a reaction-feedforward CRD $\mathcal{C}$ which stably decides whether the initial number $n=\#X$ of input $X$ is even. We withhold two copies of $X$ and let $\mathcal{C}$ stabilize on the correct output of yes. Denote $\Upsilon$ as the set of yes voters. We denote the no voters with $\overline{\Upsilon}\triangleq\Lambda\backslash\Upsilon$. Only species contained in $\Upsilon$ are present in the stable, correct output configuration.
Now, we release one of the remaining copies of $X$. We first run the chain reaction (if any) starting from only one $X$. Let $\Omega_{X}:=\{S\mid\exists\mathbf{x}\in\mathsf{reach}(\{1X\}):\mathbf{x}(S)>0\}$ be the set of species producible from $\{1X\}$ (e.g., if there is no reaction $X\mathop{\rightarrow}\limits...$, then $\Omega_{X}$ is just $\{X\}$). Without loss of generality, we assume $\Omega_{X}\subseteq\Upsilon$, that is, all of $X$’s direct products are yes-voters (if not, exchange $\Upsilon$ and $\overline{\Upsilon}$ in what follows). To correct the answer ($n+1\not\equiv n\bmod 2$), $\mathcal{C}$ must consume all species currently present and produce at least one copy of a species in $\overline{\Upsilon}$. It follows that for all $X\in\Omega_{X}$, $\mathcal{C}$ contains a reaction with $X$ as a reactant. Further, none of these reactions contain a reactant of $\overline{\Upsilon}$, since none are present in the current configuration.
Finally, we release the last remaining copy of $X$. Again, we produce the set $\Omega_{X}$ from $X$. To invert the vote again, we must consume all $Y\in\overline{\Upsilon}$ and produce at least one member of $\Upsilon$. The reaction(s) consuming $\overline{\Upsilon}$ must have a member of $\Omega_{X}$ as a reactant since the configuration is stable without $\Omega_{X}$. Further, the reaction cannot be any of the ones from before since they contain a member of $\overline{\Upsilon}$ as reactant.
Since there are least two reactions sharing a common species as reactant, the reactions cannot be ordered such that no reactant of the first of these reactions appears in the latter one. This makes $\mathcal{C}$ non-reaction-feedforward, contradicting our initial assumption.
{lemmarep}
Reaction-feedforward CRDs can’t stably decide the majority predicate $[X_{1}\geq X_{2}?]$.
Proof 6.33.
Suppose, for the sake of contradiction, there exists a reaction-feedforward CRD $\mathcal{C}$ which stably decides the predicate. We let $\mathcal{C}$ stabilize on input $\{nX_{1},nX_{2}\}$ (yielding output yes), while withholding two copies of $X_{2}$. We release one $X_{2}$. Again, we consider the full set of species a single $X_{2}$ could produce before reacting with other molecules (denoted $\Omega_{X_{2}}$). Without loss of generality, we consider all of them yes voters i.e. $\Omega_{X_{2}}\subseteq\Upsilon$. The correct output now changes to no, and all yes voters must be consumed by reactions that only have reactants which are yes voters and further, these reactions contain all species in $\Omega_{X_{2}}$ as reactants.
Once the vote has reversed and stabilized, it contains only species of $\overline{\Upsilon}\triangleq\Lambda\backslash\Upsilon$. We release the last $X_{2}$ and let it produce $\Omega_{X_{2}}$. Since $\Omega_{X_{2}}\subseteq\Upsilon$ i.e. all elements are yes voters, but the correct vote is still no, all $X\in\Omega_{X_{2}}$ must be consumed again. This time, they must be consumed in reactions involving no voters, must be distinct reactions from those in the previous step. Thus, all species $X\in\Omega_{X_{2}}$ appear at least twice as a reactant, breaking the reaction-feedforward condition.
@@ 7 Conclusion
[…]
full We explored the computational capabilities of execution bounded Chemical Reaction Networks (CRNs), which terminate after a finite number of reactions. This constraint aligns the model with practical scenarios where fuel supply is limited.
@@
A key question remains open: Can execution bounded CRNs compute semilinear functions and predicates within polylogarithmic time? Angluin, Aspnes and Eisenstat [2] introduced a fast population protocol that simulates a register machine with high probability, and can be made probability 1 with semilinear predicates. However, this construction seems inherently unbounded in executions.

### v4

@@ 6.3 Impossibility of stably deciding not eventually constant predicates
Assume WLOG that $\phi(\mathbf{x}_{0})=0$. For each $i\in\mathbb{N}$, let $\mathbf{v}_{i}=\mathbf{x}_{i+1}-\mathbf{x}_{i}$, noting by condition (LABEL:lem:semilinear-not-almost-constant-alternating-period-vector-and-offset:cond2) that $\mathbf{v}_{i}\geq 0$.
We consider the sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\mathbf{a}_{2},\dots$ defined as follows. Let $\mathbf{a}_{0}$ be a stable configuration reachable from $\mathbf{x}_{0}$; since the correct answer is no, all species present in $\mathbf{a}_{0}$ vote no. Now add $\mathbf{v}_{0}$ to $\mathbf{a}_{0}$. By additivity, the configuration $\mathbf{a}_{0}+\mathbf{v}_{0}$ is reachable from $\mathbf{x}_{1}=\mathbf{x}_{0}+\mathbf{v}_{0}$. Since the correct answer for $\mathbf{x}_{1}$ is yes, $\mathcal{D}$ must go from $\mathbf{a}_{0}+\mathbf{v}_{0}$ to a stable “yes” configuration, call this $\mathbf{a}_{1}$. Now add $\mathbf{v}_{1}$ to $\mathbf{a}_{1}$. Since the correct answer is no, $\mathcal{D}$ must now reach from $\mathbf{a}_{1}+\mathbf{v}_{1}$ to a stable “no” configuration, call it $\mathbf{a}_{2}$. By condition (LABEL:lem:semilinear-not-almost-constant-alternating-period-vector-and-offset:cond3), each $\mathbf{v}_{i}$ obeys $\|\mathbf{v}_{i}\|<c$ for some constant $c$.
@@
Since $\mathcal{D}$ is entirely execution bounded, by LABEL:thm:potential-function-iff-entirely-execution-bounded, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{w}\cdot\mathbf{x}$, where $\mathbf{w}\geq\mathbf{0}$. Adding $\mathbf{v}_{i}$ to $\mathbf{a}_{i}$ increases $\Phi$ by $\mathbf{w}(\mathbf{v}_{i})$, which is bounded above by a constant since $\|\mathbf{v}_{i}\|<c$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $\mathbf{v}_{i}$ and then going from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{w}(\mathbf{v}_{i})+2$).
@@
The statement of Theorem˜6.7 does not mention the concept of a leader, but it would typically apply to leaderless CRDs. A CRD may be execution bounded from configurations with a single leader, but not execution bounded when multiple leaders are present (preventing the use of LABEL:thm:potential-function-iff-entirely-execution-bounded, which requires the CRD to be execution bounded from all configurations). For example, in Lemma˜4.7, reaction (9) occurs finitely many times if the leader/voter $S_{Y}$ or $S_{N}$ has count 1. However, if $S_{Y}$ and $S_{N}$ can be present simultaneously (e.g., if we start with two leaders), then the reactions $S_{Y}+V_{NN}\mathop{\rightarrow}\limits S_{Y}+V_{YN}$ and $S_{N}+V_{YN}\mathop{\rightarrow}\limits S_{N}+V_{NN}$ can flip between $V_{NN}$ and $V_{YN}$ infinitely often in an unbounded execution.
@@
Lemma 6.12.
lem:leaderless-and-execution-bdd-implies-entirely-execution-bdd If a leaderless CRD or CRC is execution bounded, then it is entirely execution bounded.
Proof sketch.
Since $\mathcal{C}$ is leaderless, the sum of two valid initial configurations is also valid. Thus if we can produce some species from a valid initial configuration, we can produce arbitrarily large counts of all species by adding up sufficiently many initial configurations. This means that for any configuration $\mathbf{x}$, from any sufficiently large valid initial configuration $\mathbf{i}$, some $\mathbf{y}\geqq\mathbf{x}$ is reachable from $\mathbf{i}$. But if $\mathcal{C}$ is execution bounded from $\mathbf{i}$, since $\mathbf{i}\Rightarrow\mathbf{y}$, it must also be execution bounded from $\mathbf{y}$, thus also from $\mathbf{x}$ since by additivity any reactions applicable to $\mathbf{x}$ are also applicable to $\mathbf{y}$. ∎
Proof 6.13.
@@
LABEL:lem:leaderless-and-execution-bdd-implies-entirely-execution-bdd lets us replace “entirely execution bounded” in LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant with “leaderless and execution bounded”:
@@
In particular, since the original model of population protocols [angluin2004computation] defined them as leaderless and all-voting—and since population protocols are non-collapsing—we have the following.
@@ 7 Conclusion
We explored the computational capabilities of execution bounded Chemical Reaction Networks (CRNs), which terminate after a finite number of reactions. This constraint aligns the model with practical scenarios where fuel supply is limited.
@@
A key question remains open: Can execution bounded CRNs compute semilinear functions and predicates within polylogarithmic time? Angluin, Aspnes and Eisenstat [AngluinAE2008Fast] introduced a fast population protocol that simulates a register machine with high probability, and can be made probability 1 with semilinear predicates. However, this construction seems inherently unbounded in executions.